# TE-PAI: Exact Time Evolution by Sampling Random Circuits

### v2

@@ Abstract
Simulating time evolution under quantum Hamiltonians is one of the most natural applications of quantum computers. We introduce TE-PAI, which simulates time evolution exactly by sampling random quantum circuits for the purpose of estimating observable expectation values at the cost of an increased circuit repetition. The approach builds on the Probabilistic Angle Interpolation (PAI) technique and we prove that it simulates time evolution without discretisation or algorithmic error while achieving shallow circuit depths with optimal scaling that saturates the Lieb-Robinson bound. Another significant advantage of TE-PAI is that it only requires executing random circuits that consist of Pauli rotation gates of only two kinds of rotation angles ±$\pm\Delta$ and $\pi$, along with measurements. While TE-PAI is highly beneficial for NISQ devices, we additionally develop an optimised early fault-tolerant implementation using catalyst circuits and repeat-until-success teleportation, concluding that the approach requires orders of magnitude fewer T-states than conventional techniques, such as Trotterization – we estimate $3\times 10^{5}$ T states are sufficient for the fault-tolerant simulation of a $100$-qubit Heisenberg spin Hamiltonian. Furthermore, TE-PAI allows for a highly configurable trade-off between circuit depth and measurement overhead by adjusting the rotation angle $\Delta$ arbitrarily. We expect that the approach will be a major enabler in the late NISQ and early fault-tolerant periods as it can compensate circuit-depth and qubit-number limitations through an increased circuit repetition.
@@ I Introduction
Accurately modelling the time evolution of quantum systems is an important task but presents a significant challenge in classical computing. Thus, simulating quantum dynamics is regarded as one of the most promising applications of quantum computers [1, 2] and may provide an exponential speedup over classical computers. The simplest such approach, the Trotter-Suzuki decomposition [3, 4], approximates the time evolution through a relatively simple circuit that contains evolutions under the individual Hamiltonian terms. A drawback of the approach is that circuits may need to be quite deep to sufficiently suppress approximation errors. This discretisation error, also called the Trotter error, is inevitable with finite circuit depth and can be particularly daunting in, e.g., quantum chemistry applications [5, 6, 7, 8]. These issues are further exacerbated when the aim is to simulate dynamics under time-dependent Hamiltonians, as we demonstrate below. Indeed, sophisticated quantum algorithms, such as linear combination of unitaries (LCU) [9, 10, 11], quantum signal processing [12, 13] or quantum walks [14, 15], can achieve a fundamentally improved circuit-depth scaling compared to Trotterisation; however, they require significant overheads in quantum resources.
@@
We make significant progress and develop TE-PAI, which: a) simulates “effectively exact” time evolution on average; b) requires executing only very simple circuits and performing measurements on them, i.e., it does not require advanced quantum resources such as ancillary qubits or controlled evolutions; c) can naturally simulate time evolution under time-dependent Hamiltonians; d) achieves shallow circuit depths with optimal scaling that saturates the Lieb-Robinson bound. TE-PAI is “effectively exact” in the sense that one needs to choose a priori a desired precision $\epsilon$ to which the expectation value of the time-evolved observable is estimated, i.e., the level of statistical uncertainty. While algorithmic errors associated with time evolution are finite, they can be suppressed arbitrarily below $\epsilon$ without increasing circuit depths, i.e., by committing $O(\epsilon^{-1})$ classical pre-processing resources.
@@
TE-PAI proceeds by constructing an unbiased estimator for the entire exact time-evolution superoperator – this allows us to estimate expectation values of time-evolved observables by sampling the output of quantum circuits and is thus compatible with advanced measurement techniques, such as classical shadows [16, 17] or Pauli grouping techniques [18, 19]. This immediately enables a wide range of applications, such as shadow spectroscopy for estimating energy gaps in the problem Hamiltonian [20]. TE-PAI can be combined straightforwardly with a broad range of further techniques that require the estimation of time-dependent correlators. Specifically, one can estimate the Loschmidt echoes $\langle\psi(0)|\psi(t)\rangle$ and $\langle\psi(t)|O|\psi(t^{\prime})\rangle$ of fundamental importance, by estimating the Pauli $Z$ expectation value on an ancilla qubit. This enables the estimation of ground and excited state energies via statistical phase estimation [21, 22, 23], the estimation of the density of states [24], as well as the estimation of expectation values of observables $O$ in eigenstates of the problem Hamiltonian [25]. These techniques are expected to unlock some of the most important application areas of early fault-tolerant quantum computers, e.g., obtaining ground-state energies, correlation functions and molecular configurations, etc. However, we note that our approach is based on expected value estimation and is thus not compatible with conventional phase estimation or with other “single-shot” techniques.
@@
Previous work, such as qDRIFT [26, 27, 28], similarly use randomisation but the circuit depths are not independent of the approximation error. Prior works [30, 29] achieve some of the advantages of TE-PAI (exact evolution, comparable circuit depths), however, require more complex controlled circuits and are not compatible with advanced measurement techniques, such as classical shadows [16, 17] or Pauli grouping techniques [18, 19]. For instance,  [29] requires estimating observables one-by-one using Hadamard tests and thus requires random circuits to be controlled on an ancillary qubit, as we detail in Appendix A. In contrast, TE-PAI only requires executing simple random circuits and measuring their outputs, and thus benefits from almost unlimited compatibility with a broad range of applications and can be combined naturally with all error mitigation techniques, classical shadows, and randomised protocols that use time evolution as a subroutine, such as spectroscopy or estimating the density of states and beyond [16, 20, 24].
@@
The main technical tool we exploit is Probabilistic Angle Interpolation (PAI) [31], which is particularly relevant for near-term and early fault-tolerant applications, where averaging over many circuit executions is required for expected-value measurements. While PAI increases the sampling cost, it provably achieves the least possible overhead. TE-PAI applies this mathematical formalism to time evolution circuits, generating random circuits with only two kinds of rotation angles, ±$\pm\Delta$ and $\pi$ as illustrated in Fig. 1; post-processing their outputs yields exact time-evolved expected values.
@@
We prove that TE-PAI requires a number of gates proportional to the total simulation time $T$ and the system size, thus saturating the Lieb-Robinson bound [32, 33]. A significant advantage of TE-PAI is that it offers a trade-off between sampling overhead and circuit depth, i.e., one can use NISQ-friendly shallow circuits at the cost of an increased sampling overhead. While these features are particularly important in the NISQ era, we construct optimised implementations for early fault-tolerant quantum computers (FTQC), whereby circuit-depth and width limitations will be similarly crucial.
@@
Figure 1: A single random circuit instance of TE-PAI – by executing multiple such random circuits and post-processing their measurement outcomes, one can implement effectively exact time evolution on average via 2. In the present example, we consider a 5-qubit Hamiltonian defined in Eq. 13 and a rotation angle $\Delta=\pi/2^{6}=\pi/64$. TE-PAI then uses the Pauli gates $RXX,RYY,RZZ$ and $RZ$ only with rotation angles ±$\pm\Delta=\pm\pi/64$ and only rarely with $\pi$ (gate highlighted by dotted rectangle) – when the angle $\pi$ is chosen then all measurement outcomes are multiplied by a factor $-1$. The example considers a short time evolution of $T=0.05$ which is the reason for obtaining a shallow circuit with an expected number of gates $\nu_{\infty}\approx 25$. We note that existing compilation techniques, including ones that were specifically developed for Trotterised circuit structures [34], can be applied immediately to reduce the circuit depth.
@@ I.1.1 Time-independent Hamiltonians
Each term represents the evolution under one component of the Hamiltonian $h_{k}$ for a short time interval $\frac{T}{N}$ and the approximation becomes increasingly more accurate as $N$ increases. While in the present work we are focusing on the above first-order Trotter decomposition we note that our results can immediately be applied to higher-order Trotter decompositions [35, 36].
We now briefly summarise the error analysis for time-independent Trotter decompositions following ref. [35].
@@
$\left\|\prod_{k=1}^{L}e^{-ic_{k}h_{k}\frac{T}{N}}-e^{-iH\frac{T}{N}}\right\|\leq\frac{T^{2}}{2N^{2}}\lVert c\rVert^{2}_{T},$
where the Trotterisation error norm was defined in ref. [35] as
@@
Indeed, the additive Trotter error, $\epsilon$, can be reduced at the expense of proportionally increasing the circuit depth. In the next section we describe our algorithm which utilizes a probabilistic approach to generate random circuits from these Trotter-Suzuki circuit templates. In stark contrast, the circuit depth in our approach is independent of the precision, and $N$ is a parameter that only affects the complexity of classical pre-processing.
@@ I.1.2 Time-dependent Hamiltonians
Building on the previous time-independent case, we extend the formalism to time-dependent Hamiltonians, which are crucial for accurately simulating complex quantum systems in practice that evolve under time-dependent interactions, such as in quantum control [37]. In particular, we consider a Hamiltonian $H(t)$ whose decomposition coefficients $c_{k}(t)$ are time-dependent as
@@
We will assume that $c_{k}(t)$ are absolutely continuous functions of time. This assumption is necessary to bound the convergence rate in our proofs in Appendices D and C. We then define their average $\ell_{1}$-norm as
@@
The approximation assumes that the Hamiltonian remains constant within each small interval $\tfrac{T}{N}$, this approximation improves as $N$ is increased, and error bounds have been reported in ref. [36]. We also note that a generalisation of the above product formula can be found in ref. [36] which has been derived by truncating the Magnus expansion [38] to the first order.
@@ II.1 Unbiased estimator for general product formulas
We use the Probabilistic Angle Interpolation (PAI) technique [31], which we summarise in Appendix B, and which builds on the observation that estimating an expected value of an observable requires a quantum circuit to be run and measured many times. At each circuit run, PAI randomly replaces the angle $\theta_{kj}$ in the rotation gate $R_{k}(\theta_{kj})$ with one of only three discrete rotation angles 00, $\operatorname{sign}(\theta_{kj})\Delta$, or $\pi$. We will denote these three gate variants as
@@
The crucial observation that TE-PAI exploits is that in the limit $N\rightarrow\infty$, where the Trotter circuit approaches exact evolution, the rotation angles $\theta_{kj}$ become vanishingly small. Thus, with probability nearly equal to 1, we almost always choose the first gate variant, the identity operation (see Appendix C). As we prove below, this ultimately guarantees that the total circuit depth remains finite even in the limit $N\rightarrow\infty$ and thus the parameter $N$ only influences the complexity of classical pre-computation.
We review in detail in Appendix B that the PAI approach builds on the fact that the superoperator representation $\mathcal{R}_{k}(\theta_{kj})$ of each unitary gate $R_{k}(\theta_{kj})$ can be decomposed analytically as
@@
where $\mathcal{A}$, $\mathcal{B}_{kj}$ and $\mathcal{C}_{k}$ are superoperator representations of the unitary gates in Eq. 7 which generally act isomorphically via conjugation, e.g., $\mathcal{R}_{k}\mathrm{vec}[\rho]=\mathrm{vec}[R_{k}\rho R_{k}^{\dagger}]$. The coefficients $\gamma_{l}(\theta)$ are provided explicitly as trigonometric functions in Appendix B.
Focusing on a single gate element ^$\hat{\mathcal{R}}_{k}(\theta_{kj})$, our classical pre-processing algorithm randomly selects one of the three discrete gate variants as $\hat{\mathcal{D}}_{l}\in\{\mathcal{A},\mathcal{B}_{kj},\mathcal{C}_{k}\}$ according to the probabilities $p_{l}=|\gamma_{l}(\theta_{kj})|/\lVert\gamma(\theta_{kj})\rVert_{1}$ for $l\in\{1,2,3\}$ in order to sample the unbiased estimator of the desired, continuous-angle gate $\mathcal{R}_{k}(\theta_{kj})$ as
^$\hat{\mathcal{R}}_{k}(\theta_{kj})=\|\gamma(|\theta_{kj}|)\|_{1}\operatorname{sign}[\gamma_{l}(|\theta_{kj}|)]\hat{\mathcal{D}}_{l}.$ (9)
Ref. [31] analytically proved that this choice of the three discrete angle settings minimises the measurement overhead characterised by $\|\gamma(|\theta_{kj}|)\|_{1}$, confirming numerical optimisation results of ref. [39].
Replacing each gate in the product formula in Eq. 6 with the above unbiased estimator then allows us to construct an unbiased estimator for the entire time evolution operator. As we now summarise, this is a direct consequence of Statement 2 of ref. [31].
@@
using the unbiased estimators ^$\hat{\mathcal{R}}_{k}(\theta_{kj})$ of the individual continuous-angle rotations from Eq. 9. Specifically, the mean value of the estimator is $\mathds{E}[\hat{\mathcal{U}}]=\mathcal{U}$. The classical computational complexity of generating $N_{s}$ random circuits is $O(NLN_{s})$.
Please refer to Appendix B for a detailed derivation. Fig. 1 shows an example of a random circuit generated using the PAI approach that simulates the time evolution under the Hamiltonian Eq. 13 for 5-qubit. Finally our protocol is summarised as follows.
@@
  • Generate $N_{s}$ random circuits by randomly replacing gates in the circuit $\mathcal{U}$ with fixed rotation angles of ±$\pm\Delta$ and $\pi$ according to the PAI protocol in 2.
  • Execute all random circuits and in post-processing multiply all measurement outcomes with their corresponding prefactor $\prod_{j=1}^{N}\prod_{k=1}^{L}\|\gamma(|\theta|_{kj})\|_{1}\operatorname{sign}[\gamma_{l}(|\theta_{kj}|)]$ where the index $l:=l_{kj}$ is chosen randomly for each gate in the circuit as $l\in\{1,2,3\}$. When classical shadows are estimated then the expected value from each snapshot needs to be multiplied by this factor as detailed in ref. [16].

### v3

@@ Abstract
Simulating time evolution under quantum Hamiltonians is one of the most natural applications of quantum computers. We introduce TE-PAI, which simulates time evolution exactly by sampling random quantum circuits for the purpose of estimating observable expectation values at the cost of an increased circuit repetition. The approach builds on the Probabilistic Angle Interpolation (PAI) technique and we prove that it simulates time evolution without discretisation or algorithmic error while achieving shallow circuit depths with optimal scaling that saturates the Lieb-Robinson bound. Another significant advantage of TE-PAI is that it only requires executing random circuits that consist of Pauli rotation gates of only two kinds of rotation angles $\pm\Delta$ and $\pi$, along with measurements. While TE-PAI is highly beneficial for NISQ devices, we additionally develop an optimised early fault-tolerant implementation using catalyst circuits and repeat-until-success teleportation, concluding that the approach requires orders of magnitude fewer T-states than conventional techniques, such as Trotterization – we estimate $3\times 10^{5}$ T states are sufficient for the fault-tolerant simulation of a $100$-qubit Heisenberg spin Hamiltonian. Furthermore, TE-PAI allows for a highly configurable trade-off between circuit depth and measurement overhead by adjusting the rotation angle $\Delta$ arbitrarily. We expect that the approach will be a major enabler in the late NISQ and early fault-tolerant periods as it can compensate circuit-depth and qubit-number limitations through an increased circuit repetition.
@@ I Introduction
Accurately modelling the time evolution of quantum systems is an important task but presents a significant challenge in classical computing. Thus, simulating quantum dynamics is regarded as one of the most promising applications of quantum computers [Feynman, Lloyd] and may provide an exponential speedup over classical computers. The simplest such approach, the Trotter-Suzuki decomposition [Suzuki1, Suzuki2], approximates the time evolution through a relatively simple circuit that contains evolutions under the individual Hamiltonian terms. A drawback of the approach is that circuits may need to be quite deep to sufficiently suppress approximation errors. This discretisation error, also called the Trotter error, is inevitable with finite circuit depth and can be particularly daunting in, e.g., quantum chemistry applications [chemistry1, chemistry2, chemistry3, chemistry4]. These issues are further exacerbated when the aim is to simulate dynamics under time-dependent Hamiltonians, as we demonstrate below. Indeed, sophisticated quantum algorithms, such as linear combination of unitaries (LCU) [LCU1, LCU2, LCU3], quantum signal processing [QSP1, QSP2] or quantum walks [qw1, qw2], can achieve a fundamentally improved circuit-depth scaling compared to Trotterisation; however, they require significant overheads in quantum resources.
@@
We make significant progress and develop TE-PAI, which: a) simulates “effectively exact” time evolution on average; b) requires executing only very simple circuits and performing measurements on them, i.e., it does not require advanced quantum resources such as ancillary qubits or controlled evolutions; c) can naturally simulate time evolution under time-dependent Hamiltonians; d) achieves shallow circuit depths with optimal scaling that saturates the Lieb-Robinson bound. TE-PAI is “effectively exact” in the sense that one needs to choose a priori a desired precision $\epsilon$ to which the expectation value of the time-evolved observable is estimated, i.e., the level of statistical uncertainty (see LABEL:app:classical_cost for details). While algorithmic errors associated with time evolution are finite, they can be suppressed arbitrarily below $\epsilon$ without increasing circuit depths, i.e., by committing $O(\epsilon^{-1})$ classical pre-processing resources.
@@
TE-PAI proceeds by constructing an unbiased estimator for the entire exact time-evolution superoperator – this allows us to estimate expectation values of time-evolved observables by sampling the output of quantum circuits and is thus compatible with advanced measurement techniques, such as classical shadows [mitigation_shadow, shadow] or Pauli grouping techniques [Crawford2019, jena2019pauli]. This immediately enables a wide range of applications, such as shadow spectroscopy for estimating energy gaps in the problem Hamiltonian [chan2022algorithmic]. TE-PAI can be combined straightforwardly with a broad range of further techniques that require the estimation of time-dependent correlators. Specifically, one can estimate the Loschmidt echoes $\langle\psi(0)|\psi(t)\rangle$ and $\langle\psi(t)|O|\psi(t^{\prime})\rangle$ of fundamental importance, by estimating the Pauli $Z$ expectation value on an ancilla qubit. This enables the estimation of ground and excited state energies via statistical phase estimation [Wang_2023, Lin_2022, Ding2023simultaneous], the estimation of the density of states [goh2024direct], as well as the estimation of expectation values of observables $O$ in eigenstates of the problem Hamiltonian [PRXQuantum.2.040361]. These techniques are expected to unlock some of the most important application areas of early fault-tolerant quantum computers, e.g., obtaining ground-state energies, correlation functions and molecular configurations, etc. However, we note that our approach is based on expected value estimation and is thus not compatible with conventional phase estimation or with other “single-shot” techniques.
@@
Previous work, such as qDRIFT [qDRIFT1, qDRIFT2, qDRIFT3], similarly use randomisation but the circuit depths are not independent of the approximation error. Prior works [dyson-base, similar] achieve some of the advantages of TE-PAI (exact evolution, comparable circuit depths), however, require more complex controlled circuits and are not compatible with advanced measurement techniques, such as classical shadows [mitigation_shadow, shadow] or Pauli grouping techniques [Crawford2019, jena2019pauli]. For instance,  [similar] requires estimating observables one-by-one using Hadamard tests and thus requires random circuits to be controlled on an ancillary qubit, as we detail in Appendix A. In contrast, TE-PAI only requires executing simple random circuits and measuring their outputs, and thus benefits from almost unlimited compatibility with a broad range of applications and can be combined naturally with all error mitigation techniques, classical shadows, and randomised protocols that use time evolution as a subroutine, such as spectroscopy or estimating the density of states and beyond [mitigation_shadow, chan2022algorithmic, goh2024direct].
@@
The main technical tool we exploit is Probabilistic Angle Interpolation (PAI) [PAI], which is particularly relevant for near-term and early fault-tolerant applications, where averaging over many circuit executions is required for expected-value measurements. While PAI increases the sampling cost, it provably achieves the least possible overhead. TE-PAI applies this mathematical formalism to time evolution circuits, generating random circuits with only two kinds of rotation angles, $\pm\Delta$ and $\pi$ as illustrated in Fig. 1; post-processing their outputs yields exact time-evolved expected values. Conventional trotterisation requires small gate angle settings that are proportional to the individual Hamiltonian terms. However, in real experiments the gate angles always have a finite precision as well as added coherent over or under rotations, which effectively leads to implementing the correct trotter evolution of a different Hamiltonian with slightly different Pauli coefficients – the effect of this may be radical near phase transitions. Our randomised approach effectively overcomes this limitation as the rotation angle $\Delta$ can be chosen arbitrarily in our approach and can be updated according to our best knowledge of the true, experimentally calibrated rotation angle.
@@
We prove that TE-PAI requires a number of gates proportional to the total simulation time $T$ and the system size, thus saturating the Lieb-Robinson bound [LR, Haah2021]. A significant advantage of TE-PAI is that it offers a trade-off between sampling overhead and circuit depth, i.e., one can use NISQ-friendly shallow circuits at the cost of an increased sampling overhead. While these features are particularly important in the NISQ era, we construct optimised implementations for early fault-tolerant quantum computers (FTQC), whereby circuit-depth and width limitations will be similarly crucial.
@@
Figure 1: A single random circuit instance of TE-PAI – by executing multiple such random circuits and post-processing their measurement outcomes, one can implement effectively exact time evolution on average via 2. In the present example, we consider a 5-qubit Hamiltonian defined in Eq. 13 and a rotation angle $\Delta=\pi/2^{6}=\pi/64$. TE-PAI then uses the Pauli gates $RXX,RYY,RZZ$ and RZRZ only with rotation angles $\pm\Delta=\pm\pi/64$ and only rarely with $\pi$ (gate highlighted by dotted rectangle) – when the angle $\pi$ is chosen then all measurement outcomes are multiplied by a factor 1-1. The example considers a short time evolution of $T=0.05$ which is the reason for obtaining a shallow circuit with an expected number of gates $\nu_{\infty}\approx 25$. We note that existing compilation techniques, including ones that were specifically developed for Trotterised circuit structures [2qan], can be applied immediately to reduce the circuit depth.
@@ I.1.1 Time-independent Hamiltonians
Each term represents the evolution under one component of the Hamiltonian $h_{k}$ for a short time interval $\frac{T}{N}$ and the approximation becomes increasingly more accurate as $N$ increases. While in the present work we are focusing on the above first-order Trotter decomposition we note that our results can immediately be applied to higher-order Trotter decompositions [trotter_err, oscillate].
We now briefly summarise the error analysis for time-independent Trotter decompositions following ref. [trotter_err].
@@
$\epsilon_{T}:=\left\|\left(\prod_{k=1}^{L}e^{-ic_{k}h_{k}\frac{T}{N}}\right)^{N}-e^{-iHT}\right\|\leq\frac{T^{2}}{2N}\lVert c\rVert^{2}_{T},$
where the Trotterisation error norm was defined in ref. [trotter_err] as
@@
Indeed, the additive Trotter error, $\epsilon$, can be reduced at the expense of proportionally increasing the circuit depth. In the next section we describe our algorithm which utilizes a probabilistic approach to generate random circuits from these Trotter-Suzuki circuit templates. In stark contrast, the circuit depth in our approach is independent of the precision, and $N$ is a parameter that only affects the complexity of classical pre-processing as we detail in LABEL:app:classical_cost.
@@ I.1.2 Time-dependent Hamiltonians
Building on the previous time-independent case, we extend the formalism to time-dependent Hamiltonians, which are crucial for accurately simulating complex quantum systems in practice that evolve under time-dependent interactions, such as in quantum control [koch2022quantum]. In particular, we consider a Hamiltonian $H(t)$ whose decomposition coefficients $c_{k}(t)$ are time-dependent as
@@
We will assume that $c_{k}(t)$ are absolutely continuous functions of time. This assumption is necessary to bound the convergence rate in our proofs in LABEL:app:proof1 and LABEL:app:proof2. We then define their average $\ell_{1}$-norm as
@@
The approximation assumes that the Hamiltonian remains constant within each small interval $\tfrac{T}{N}$, this approximation improves as $N$ is increased, and error bounds have been reported in ref. [oscillate]. We also note that a generalisation of the above product formula can be found in ref. [oscillate] which has been derived by truncating the Magnus expansion [magnus] to the first order.
@@ II.1 Unbiased estimator for general product formulas
We use the Probabilistic Angle Interpolation (PAI) technique [PAI], which we summarise in LABEL:app:PAI, and which builds on the observation that estimating an expected value of an observable requires a quantum circuit to be run and measured many times. At each circuit run, PAI randomly replaces the angle $\theta_{kj}$ in the rotation gate $R_{k}(\theta_{kj})$ with one of only three discrete rotation angles 0, $\operatorname{sign}(\theta_{kj})\Delta$, or $\pi$. We will denote these three gate variants as
@@
The crucial observation that TE-PAI exploits is that in the limit $N\rightarrow\infty$, where the Trotter circuit approaches exact evolution, the rotation angles $\theta_{kj}$ become vanishingly small. Thus, with probability nearly equal to 1, we almost always choose the first gate variant, the identity operation (see Appendix LABEL:app:proof2). As we prove below, this ultimately guarantees that the total circuit depth remains finite even in the limit $N\rightarrow\infty$ and thus the parameter $N$ only influences the complexity of classical pre-computation.
We review in detail in LABEL:app:PAI that the PAI approach builds on the fact that the superoperator representation $\mathcal{R}_{k}(\theta_{kj})$ of each unitary gate $R_{k}(\theta_{kj})$ can be decomposed analytically as
@@
where $\mathcal{A}$, $\mathcal{B}_{kj}$ and $\mathcal{C}_{k}$ are superoperator representations of the unitary gates in Eq. 7 which generally act isomorphically via conjugation, e.g., $\mathcal{R}_{k}\mathrm{vec}[\rho]=\mathrm{vec}[R_{k}\rho R_{k}^{\dagger}]$. The coefficients $\gamma_{l}(\theta)$ are provided explicitly as trigonometric functions in LABEL:app:PAI.
Focusing on a single gate element $\hat{\mathcal{R}}_{k}(\theta_{kj})$, our classical pre-processing algorithm randomly selects one of the three discrete gate variants as $\hat{\mathcal{D}}_{l}\in\{\mathcal{A},\mathcal{B}_{kj},\mathcal{C}_{k}\}$ according to the probabilities $p_{l}=|\gamma_{l}(\theta_{kj})|/\lVert\gamma(\theta_{kj})\rVert_{1}$ for $l\in\{1,2,3\}$ in order to sample the unbiased estimator of the desired, continuous-angle gate $\mathcal{R}_{k}(\theta_{kj})$ as
$\hat{\mathcal{R}}_{k}(\theta_{kj})=\|\gamma(|\theta_{kj}|)\|_{1}\operatorname{sign}[\gamma_{l}(|\theta_{kj}|)]\hat{\mathcal{D}}_{l}.$ (9)
Ref. [PAI] analytically proved that this choice of the three discrete angle settings minimises the measurement overhead characterised by $\|\gamma(|\theta_{kj}|)\|_{1}$, confirming numerical optimisation results of ref. [koczor2024sparse].
Replacing each gate in the product formula in Eq. 6 with the above unbiased estimator then allows us to construct an unbiased estimator for the entire time evolution operator. As we now summarise, this is a direct consequence of Statement 2 of ref. [PAI].
@@
using the unbiased estimators $\hat{\mathcal{R}}_{k}(\theta_{kj})$ of the individual continuous-angle rotations from Eq. 9. Specifically, the mean value of the estimator is $\mathds{E}[\hat{\mathcal{U}}]=\mathcal{U}$. The classical computational complexity of generating $N_{s}$ random circuits is $O(NLN_{s})$.
Please refer to LABEL:app:PAI for a detailed derivation. Fig. 1 shows an example of a random circuit generated using the PAI approach that simulates the time evolution under the Hamiltonian Eq. 13 for 5-qubit. Finally our protocol is summarised as follows.
@@
  • Generate $N_{s}$ random circuits by randomly replacing gates in the circuit $\mathcal{U}$ with fixed rotation angles of $\pm\Delta$ and $\pi$ according to the PAI protocol in 2.
  • Execute all random circuits and in post-processing multiply all measurement outcomes with their corresponding prefactor $\prod_{j=1}^{N}\prod_{k=1}^{L}\|\gamma(|\theta|_{kj})\|_{1}\operatorname{sign}[\gamma_{l}(|\theta_{kj}|)]$ where the index $l:=l_{kj}$ is chosen randomly for each gate in the circuit as $l\in\{1,2,3\}$. When classical shadows are estimated then the expected value from each snapshot needs to be multiplied by this factor as detailed in ref. [mitigation_shadow].